OpenStudy (xxstudyingstuffxx):

Which three ordered pairs complete the table to give solutions of the equation y = x – 5? A. (–3, –2), (0, –5), and (3, 2) B. (–3, –2), (0, –5), and (3, –2) C. (–3, –8), (0, –5), and (3, 2) D. (–3, –8), (0, –5), and (3, –2)

OpenStudy (xxstudyingstuffxx):

@Aveline @ijlal @UliNK

OpenStudy (ulink):

I don't know this one, sorry.

OpenStudy (aveline):

Let's start out with (3, 2) y=x-5 2=3-5 Does (3,2) work)?

OpenStudy (ijlal):

you can compare your points if you compare your points on the graph that i have attached you will see D is the answer as all three ordered pairs lie on the graph solutions means points the lie on the graph hope you understand :)

OpenStudy (xxstudyingstuffxx):

Isn't D correct?